H 0 : = 84

H a : 84

= 87, = 10, n = 35, = 0.05


1
Expert's answer
2022-05-18T09:37:06-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=84"

"H_1:\mu\not=84"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|> 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{87-84}{10/\sqrt{35}}\approx1.7748"

Since it is observed that "z=1.7748<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(z>1.7748)=0.075931," and since "p=0.075931>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 84, at the "\alpha = 0.05" significance level.